# A Survey of Behavior Foundation Model: Next-Generation Whole-Body Control System of Humanoid Robots

### v3

@@ 3.1.3 Forward-backward Representation Learning
The estimation holds due to the optimal discriminator satisfies $D^{*}=\frac{p_{\mathcal{E}}}{p_{\mathcal{E}}+p_{\pi}}$ [100]. Finally, the divergence term can be estimated by training another critic network via off-policy TD learning, and the actor loss for FB-CPR is rewritten as
@@
$L_{\rm FB-CPR}=$ $-\mathbb{E}_{\bm{z}\sim\nu,\bm{s}\sim\mathcal{D}_{\rm online},\bm{a}\sim\pi_{\bm{z}}(\cdot|\bm{s})}\left[F(\bm{s},\bm{a},\bm{z})^{\top}\bm{z}\right]$ (16)
$+\alpha Q^{\rm div}(\bm{s},\bm{a},\bm{z}),$
@@
It is natural to find that FB-CPR is not a rigorous unsupervised method, as it leverages unlabeled demo data to assist motion prior learning. By aligning unsupervised RL with human-like behavioral priors from unlabeled data, FB-CPR enhances policy diversity and dataset coverage, enabling the agent to learn a rich latent space of behaviors (e.g., walking, jumping, handstands) and achieve robust zero-shot performance across diverse tasks. Experimental results demonstrate that Motivo achieves 83% success rate in motion tracking tasks, and performs 61% of the top-line performance in reward optimization tasks, surpassing DIFFUSER [101] in computational efficiency by requiring only 12 seconds per 300-step episode. Additionally, it outperforms ASE and CALM in motion diversity, achieving a score of 4.70 ($\pm$0.66), reflecting its ability to capture a broader range of behaviors.
3.2 Adaptation
Equipped with the derived BFMs through the aforementioned pre-training frameworks, we further introduce recent advancements in the adaptation techniques of BFMs. These work can be broadly categorized into two types: fine-tuning and towards hierarchical control.
3.2.1 Fine-tuning
@@
Fine-tuning of BFMs seeks to bridge the gap between general-purpose motion priors and task-specific requirements. While pre-trained BFMs capture broad motion distributions, they often lack precision for specialized tasks or novel environments. For example, [47] introduces fast adaptation techniques like residual latent adaptation (ReLA) and lookahead latent adaptation (LoLA), which allow BFMs to achieve zero-shot adaptation by using minimal online interactions after pre-training. These methods improve task performance by up to 40%, demonstrating the effectiveness of fast adaptation strategies that enable efficient task switching without requiring extensive retraining. In contrast, [48] introduces TaskTokens, which enhance goal-conditioned BFMs by generating task-specific tokens through a task encoder, enabling the BFM to perform complex tasks like motion tracking and goal-reaching with high success rates. Their method shows up to 99.75% success in tasks such as the long jump. Additionally, [49] addresses the limitation of FB representations by introducing belief-FB, which uses a transformer encoder to infer environmental dynamics, improving adaptation to unseen changes. Their approach achieves up to 2$\times$ improvement in performance under dynamic variations, demonstrating enhanced zero-shot capabilities. Future work could investigate additional post-training techniques, such as test-time scaling [102] or RL from human feedback [103], to further improve the adaptability and efficiency of BFMs and ensure alignment with human preferences in real-world applications.
3.2.2 Towards Hierarchical Control
@@
While fine-tuning techniques enhance the performance of BFMs for specific tasks through minimal modifications during test time, several pioneering works have attempted to establish a hierarchical control architecture based on BFM, which decouples high-level planning from low-level motion execution to achieve more scalable and flexible control [87, 88]. For example, UniHSI [84] introduces a unified framework for human-scene interaction by using language commands to guide a chain-of-contacts, which represents the sequence of human-object contact pairs. The system translates language inputs into structured task plans, which are then executed by a unified controller based on the AMP architecture. This framework achieves semantic alignment between language commands and physical motions, supporting diverse interactions with single or multiple objects. TokenHSI [83] further extends this line by proposing a transformer-based unified policy that tokenizes human proprioception and task states. By separating shared motor knowledge (proprioception token) from task-specific parameters (task tokens), it enables seamless multi-skill unification and flexible adaptation to novel tasks, such as skill composition (e.g., carrying while sitting), object or terrain shape variation, and long-horizon task completion.
@@
In contrast, CLoSD [86] introduces a text-driven RL controller that combines motion diffusion models with physics-based simulations for robust multi-task human character control. By utilizing a real-time diffusion planner and a motion tracking controller in a closed-loop feedback system, CLoSD can handle complex tasks like goal-reaching, striking, and human-object interactions, all controlled through text prompts and target locations. Similarly, UniPhys [85] introduces a diffusion-based behavior cloning framework unifying planning and control with diffusion forcing to handle prediction errors, enabling flexible control via text, velocity, and goal guidance for applications like dynamic obstacle avoidance and long-horizon planning. These applications demonstrate that BFMs serve as a pivotal bridge between high-level semantic instructions and low-level physical execution, leveraging pre-trained behavioral priors to enable zero-shot adaptation, multi-task generalization, and physics-aware motion synthesis. By integrating various control paradigms (e.g., transformer-based tokenization, and closed-loop diffusion planning), these highlight BFMs’ potential to democratize humanoid control across complex, real-world scenarios, from interactive robotics to dynamic environment adaptation.
4 Applications and Limitations
@@
BFMs are foreseen to significantly enhance humanoid robotics by providing a universal pre-trained controller capable of generalizing across diverse tasks. In this section, we explore the potential applications of BFMs in diverse industries macroscopically, such as healthcare robotics and gaming. Furthermore, we identify the key limitations of current BFMs, including the Sim2Real gap, data bottleneck, and embodiment generalization. Our analysis is inspired by the current development and applications of other foundation models, such as LLMs [104] and large vision models [105, 106].
@@
4.1 Applications
4.1.1 General Accelerator for Humanoid Robotics
@@
BFMs will act as a transformative general accelerator for humanoid robotics, accelerating the development and deployment of advanced WBC systems. Unlike the previous pipelines that require resource-intensive and task-specific training, BFMs eliminate training from scratch by pre-training on vast and diverse behavior datasets, embedding a rich behavioral prior that accelerates downstream adaptation exponentially [47]. Moreover, advanced BFMs with zero-shot adaptation capabilities, such as Motivo [46], can directly map high-level task specifications (e.g., goal states and reward functions) to low-level control actions, bypassing traditional RL loops entirely. This capability efficiently solves basic control tasks but also facilitates rapid prototyping, allowing developers to evaluate robot behaviors in both simulation and real-world environments within minutes, dramatically shortening the development cycle.
4.1.2 Virtual Agents and Gaming
@@
Generative AI has significantly revolutionized digital content creation, particularly in art, animation, and game design [107, 108, 109, 110, 111, 112, 113]. Among these, controlling dynamic and interactive behaviors for non-player characters (NPCs) in games remains a significant challenge. Traditional rule-based or scripted NPCs often exhibit limited diversity, unnatural movements, and poor adaptability to player actions [114, 115, 116]. BFMs offer a groundbreaking solution by enabling lifelike, context-aware NPC behaviors without extensive manual scripting. Pre-trained on diverse human behavior datasets, BFMs generate adaptive actions, such as tactical combat, social engagement, or exploration, seamlessly responding to dynamic player inputs. By integrating BFMs with LLMs, NPCs can interpret complex player instructions (e.g., dialogue-driven commands in role-playing games) and foster immersive and responsive interactions. This capability positions BFMs as a pivotal technology for revolutionizing virtual agents, enabling next-generation gaming experiences with unprecedented behavioral realism and interactivity.
4.1.3 Towards Industry 5.0
@@
While Industry 4.0 introduced smart factories with cyber-physical systems, IoT, and AI-driven automation, Industry 5.0 shifts toward human-centric, resilient, and sustainable manufacturing, emphasizing collaborative robotics, adaptive intelligence, and personalized production [117, 118, 119, 120, 121]. To that end, robots must move beyond rigid automation and instead exhibit generalizable, adaptive, and explainable behaviors [122, 123, 124, 125, 126]. BFMs are expected to empower this landscape by enabling humanoid robots to seamlessly blend pre-trained motor skills with real-time adaptability, effortlessly switching between tasks like precision welding and adaptive part handling. By integrating large multimodal models with BFMs, robots can process diverse inputs like gestures, voice commands (e.g., ”handle gently”), or environmental cues, fostering intuitive human-robot collaboration in shared workspaces. BFMs also ensure resilience, autonomously recovering from disturbances like unbalanced loads in logistics, and support personalized production through zero-shot or few-shot learning.
4.1.4 Healthcare and Assistive Robotics
@@
The global population aging presents unprecedented challenges for healthcare systems [127, 128, 129, 130], increasing demand for assistive technologies that support independent living and rehabilitation. Extensive and diverse robots have been developed for robot-assisted dressing, rehabilitation therapy, medical treatment, and caregiving for the elderly and children [131, 132, 133, 134, 135, 136, 137, 138, 139]. Humanoid robots are ideally suited for these tasks due to their anthropomorphic design, navigating human-centric environments and perform precise, natural, and intuitive interactions. BFMs offer a promising solution by enabling robots to adapt to diverse user needs and unstructured environments. For instance, BFMs can empower assistive robots to perform tasks like mobility support (e.g., fall prevention, gait assistance) or daily tasks (e.g., object retrieval, meal preparation) with minimal user-specific tuning. In rehabilitation, BFMs trained on clinician-guided demonstrations can personalize therapy protocols by dynamically adjusting task difficulty or providing real-time feedback based on patient progress.
4.2 Limitations
4.2.1 Sim2Real Gap
@@
The Sim2Real gap is a persistent challenge in robotics, representing the performance discrepancy between the policies trained in simulators and their real-world deployment [35, 36, 37, 38]. Traditional model-based controllers address this through explicit physics modeling and robust optimization, while data-driven approaches employ domain randomization and system identification [140, 141, 142, 143]. Recent advancements like ASAP [144] mitigate dynamics mismatch via residual action learning yet face policy-specific limitations, as their residuals are trained on trajectories from a single pre-trained policy, restricting generalization. BFMs exacerbate this challenge by encoding a vast spectrum of behaviors, from locomotion to multi-contact interactions, introducing high-dimensional transfer risks. For example, a BFM trained for diverse humanoid motions may fail to adapt to real-world actuator delays, friction variations, or sensor noise, resulting in unstable or unsafe execution. The integration of visual signals into control systems further compounds the challenge, and perceptual domain shifts (e.g., lighting, texture, or camera calibration mismatches) and generalization gaps in visual features can destabilize motion policies based on simulated visual inputs. Current BFMs remain largely confined to simulation, with no documented large-scale real-world deployments. While several pioneering works have been devoted to developing BFM-like controllers in real humanoid robots, like HugWBC [60] and CLONE [145], their motion skills remain narrow, highlighting a significant challenge to Sim2Real feasibility for maintaining behavior richness. This gap stems from behavioral overgeneralization, dynamics mismatches, and latent space instability, which hinder scaling targeted Sim2Real successes (e.g., quadruped locomotion or grasping) to BFM’s complexity.
4.2.2 Data Bottleneck

### v4

@@ III-A3 Forward-backward Representation Learning
The estimation holds due to the optimal discriminator satisfies $D^{*}=\frac{p_{\mathcal{E}}}{p_{\mathcal{E}}+p_{\pi}}$ [goodfellow2014@generative]. Finally, the divergence term can be estimated by training another critic network via off-policy TD learning, and the actor loss for FB-CPR is rewritten as
@@
$L_{\rm FB-CPR}=$ $-\mathbb{E}_{\begin{subarray}{c}\bm{z}\sim\nu,\\ \bm{s}\sim\mathcal{D}_{\rm online},\bm{a}\sim\pi_{\bm{z}}(\cdot|\bm{s})\end{subarray}}\left[F(\bm{s},\bm{a},\bm{z})^{\top}\bm{z}\right]$ (16)
$+\alpha Q_{r^{\rm div}}^{\pi}(\bm{s},\bm{a},\bm{z}).$
@@
It is natural to find that FB-CPR is not a rigorous unsupervised method, as it leverages unlabeled demo data to assist motion prior learning. By aligning unsupervised RL with human-like behavioral priors from unlabeled data, FB-CPR enhances policy diversity and dataset coverage, enabling the agent to learn a rich latent space of behaviors (e.g., walking, jumping, handstands) and achieve robust zero-shot performance across diverse tasks. Experimental results demonstrate that Motivo achieves an 83% success rate in motion tracking tasks, and performs 61% of the top-line performance in reward optimization tasks, surpassing DIFFUSER [janner2022planning] in computational efficiency by requiring only 12 seconds per 300-step episode. Additionally, it outperforms ASE and CALM in motion diversity, achieving a score of 4.70 ($\pm$0.66), reflecting its ability to capture a broader range of behaviors.
Moving beyond simulated environments, BFM-Zero [li2025bfm] first successfully deploys BFMs derived from the FB framework in real-world humanoid robots by introducing key architectural innovations that enhance sim-to-real robustness. Building upon the FB-CPR foundation, BFM-Zero incorporates a dual-critic architecture that augments the adversarial imitation objective with a separate critic network explicitly trained to maximize auxiliary stability rewards, such as action smoothness and joint limit avoidance. Meanwhile, BFM-zero integrates extensive domain randomization, history-based policies, and a scaled-up transformer-inspired residual network. These efforts collectively enable the learned policies to exhibit real-time, stable, and naturalistic WBC directly transferred from simulation to hardware, such as the Unitree G1 and Booster T1 robots, without any fine-tuning.
III-B Adaptation
Equipped with the derived BFMs through the aforementioned pre-training frameworks, we next introduce recent advancements in the adaptation techniques of BFMs. These approaches can be broadly categorized into two types: fine-tuning and towards hierarchical control. It is worthwhile to highlight the distinction between the adaptation capabilities arising directly from pre-training and the fine-tuning approaches presented in this section. Pre-trained BFMs often exhibit representation-level generalization, allowing them to solve downstream tasks directly through task conditioning, such as providing goal states, latent vectors, or reward embeddings, without updating model parameters. For example, Motivo [tirinzoni2025zeroshot] can generate desirable locomotion simply by constructing a reward function that specifies the intended behavior. In contrast, the adaptation techniques reviewed below often involve refining parameters or introducing additional modules to further enhance performance in specific tasks or environments.
III-B1 Fine-tuning
@@
Fine-tuning of BFMs seeks to bridge the gap between general-purpose behavioral priors and task-specific requirements. While pre-trained BFMs capture broad behavioral priors, they often lack precision for specialized tasks or novel environments. To address this, fast adaptation techniques such as residual latent adaptation (ReLA) and lookahead latent adaptation (LoLA) [sikchi2025fast] enable BFMs to achieve zero-shot adaptation with minimal online interactions after pre-training. These methods improve task performance by up to 40%, demonstrating the effectiveness of fast adaptation strategies that enable efficient task switching without requiring extensive retraining. In contrast, methods like TaskTokens [vainshtein2025task] enhance goal-conditioned BFMs by generating task-specific tokens through a task encoder, achieving high success rates across complex tasks such as motion tracking and goal-reaching. TaskTokens shows up to 99.75% success in tasks such as the long jump. Furthermore, the belief-FB method [bobrin2025zero] extends forward–backward representations by incorporating a transformer-based dynamics inference module, substantially improving robustness to unseen environmental variations and achieving a performance improvement of up to $2\times$. Future work could investigate additional post-training techniques, such as test-time scaling [gandhi2024stream] or RL from human feedback [ouyang2022training], to further improve the adaptability and efficiency of BFMs and ensure alignment with human preferences in real-world applications.
III-B2 Towards Hierarchical Control
@@
While fine-tuning techniques enhance the performance of BFMs for specific tasks through minimal modifications during test time, several pioneering works have attempted to establish a hierarchical control architecture based on BFM, which decouples high-level planning from low-level motion execution to achieve more scalable and flexible control [shao2025langwbc, xue2025leverb]. For example, UniHSI [xiao2024unified] introduces a unified framework for human-scene interaction by using language commands to guide a chain-of-contacts, which represents the sequence of human-object contact pairs. The system translates language inputs into structured task plans, which are then executed by a unified controller based on the AMP architecture. This framework achieves semantic alignment between language commands and physical motions, supporting diverse interactions with single or multiple objects. TokenHSI [pan2025tokenhsi] further extends this line by proposing a transformer-based unified policy that tokenizes human proprioception and task states. By separating shared motor knowledge (proprioception token) from task-specific parameters (task tokens), it enables seamless multi-skill unification and flexible adaptation to novel tasks, such as skill composition (e.g., carrying while sitting), variations in object or terrain shape, and long-horizon task completion.
@@
In contrast, CLoSD [tevet2025closing] introduces a text-driven RL controller that combines motion diffusion models with physics-based simulations for robust multi-task control of human characters. By utilizing a real-time diffusion planner and a motion tracking controller in a closed-loop feedback system, CLoSD can handle complex tasks, such as goal-reaching, striking, and human-object interactions, all of which are controlled through text prompts and target locations. Similarly, UniPhys [wu2025uniphys] introduces a diffusion-based behavior cloning framework that unifies planning and control with diffusion forcing to handle prediction errors, enabling flexible control via text, velocity, and goal guidance for applications, such as dynamic obstacle avoidance and long-horizon planning. These applications demonstrate that BFMs serve as a pivotal bridge between high-level semantic instructions and low-level physical execution, leveraging pre-trained behavioral priors to enable zero-shot adaptation, multi-task generalization, and physics-aware motion synthesis. By integrating various control paradigms (e.g., transformer-based tokenization and closed-loop diffusion planning), these highlight the potential of BFMs to democratize humanoid control across complex, real-world scenarios, ranging from interactive robotics to dynamic environment adaptation.
IV Applications and Limitations
@@
BFMs are foreseen to significantly enhance humanoid robotics by providing a universal pre-trained controller capable of generalizing across diverse tasks. In this section, we explore the potential applications of BFMs in various industries, including healthcare robotics and gaming. Furthermore, we identify the key limitations of current BFMs, including the Sim2Real gap, data bottleneck, and generalization to embodiment. An overview of the applications and limitations is illustrated in Fig. 9. Our analysis is inspired by the current development and applications of other foundation models, such as LLMs [naveed2025comprehensive] and large vision models [liu2024sora, awais2025foundation].
@@
IV-A Applications
IV-A1 General Accelerator for Humanoid Robotics
@@
BFMs will act as a transformative general accelerator for humanoid robotics, accelerating the development and deployment of advanced WBC systems. Unlike the previous pipelines that require resource-intensive and task-specific training, BFMs eliminate the need for training from scratch by pre-training on vast and diverse behavior datasets, thereby embedding a rich behavioral prior that accelerates downstream adaptation exponentially [sikchi2025fast]. Moreover, advanced BFMs with zero-shot adaptation capabilities, such as Motivo [tirinzoni2025zeroshot], can directly map high-level task specifications (e.g., goal states and reward functions) to low-level control actions, bypassing traditional RL loops entirely. This capability efficiently solves basic control tasks but also facilitates rapid prototyping, allowing developers to evaluate robot behaviors in both simulation and real-world environments within minutes, dramatically shortening the development cycle.
IV-A2 Virtual Agents and Gaming
@@
Virtual agents play a crucial role in advancing the development of BFMs for humanoid robots. Many foundational BFM techniques (e.g., large-scale motion pre-training and latent skill learning) are first developed and validated in simulation environments for virtual agents before being adapted for real-world humanoids. On the other hand, the strong growth of the game industry and metaverse platforms creates an increasing demand for natural, physically grounded avatar motions, which align closely with the objectives of humanoid BFMs [kopel2018implementing, mehta2022exploring, uludaugli2023non]. BFMs offer a groundbreaking solution by enabling lifelike, context-aware non-player characters (NPCs) behaviors without extensive manual scripting. Pre-trained on diverse human behavior datasets, BFMs generate adaptive actions, such as tactical combat, social engagement, or exploration, seamlessly responding to dynamic player inputs. By integrating BFMs with LLMs, NPCs can interpret complex player instructions (e.g., dialogue-driven commands in role-playing games) and foster immersive and responsive interactions. This capability positions BFMs as a pivotal technology for revolutionizing virtual agents, enabling next-generation gaming experiences with unprecedented behavioral realism and interactivity.
IV-A3 Towards Industry 5.0
@@
While Industry 4.0 introduced smart factories with cyber-physical systems, IoT, and AI-driven automation, Industry 5.0 shifts toward human-centric, resilient, and sustainable manufacturing, emphasizing collaborative robotics, adaptive intelligence, and personalized production [xu2021industry, leng2022industry, huang2022industry, akundi2022state, hassan2024systematic]. To that end, robots must move beyond rigid automation and instead exhibit generalizable, adaptive, and explainable behaviors [dzedzickis2021advanced, bartovs2021overview, arents2022smart, soori2024intelligent, lee2022does]. BFMs are expected to empower this landscape by enabling humanoid robots to seamlessly blend pre-trained motor skills with real-time adaptability, effortlessly switching between tasks like precision welding and adaptive part handling. By integrating large multimodal models with BFMs, robots can process diverse inputs like gestures, voice commands (e.g., ”handle gently”), or environmental cues, fostering intuitive human-robot collaboration in shared workspaces. BFMs also ensure resilience, autonomously recovering from disturbances like unbalanced loads in logistics, and support personalized production through zero-shot or few-shot learning.
IV-A4 Healthcare and Assistive Robotics
@@
The global population aging presents unprecedented challenges for healthcare systems [rowland2009global, bloom2015global, navaneetham2023global, zhang2025impact], resulting in an increased demand for assistive technologies that support independent living and rehabilitation. Extensive and diverse robots have been developed for robot-assisted dressing, rehabilitation therapy, medical treatment, and caregiving for the elderly and children [feil2005defining, miller2006assistive, okamura2010medical, riek2017healthcare, holland2021service, kyrarini2021survey, sanchez2021textile, zhang2022learning, javaid2025utilization]. Humanoid robots are ideally suited for these tasks due to their anthropomorphic design, navigating human-centric environments and perform precise, natural, and intuitive interactions. BFMs offer a promising solution by enabling robots to adapt to diverse user needs and unstructured environments. For instance, BFMs can empower assistive robots to perform tasks like mobility support (e.g., fall prevention, gait assistance) or daily tasks (e.g., object retrieval, meal preparation) with minimal user-specific tuning. In rehabilitation, BFMs trained on clinician-guided demonstrations can personalize therapy protocols by dynamically adjusting task difficulty or providing real-time feedback based on patient progress.
IV-B Limitations
IV-B1 Sim2Real Gap
@@
The Sim2Real gap is a persistent challenge in robotics, representing the performance discrepancy between the policies trained in simulators and their real-world deployment [kadian2020sim2real, hofer2021sim2real, iyengar2023sim2real, su2024sim2real]. Traditional model-based controllers address this through explicit physics modeling and robust optimization, while data-driven approaches employ domain randomization and system identification [kaspar2020sim2real, horvath2022object, huber2024domain, yao2025sim2real]. Recent advancements like ASAP [he2025asap] mitigate dynamics mismatch via residual action learning yet face policy-specific limitations, as their residuals are trained on trajectories from a single pre-trained policy, thereby restricting generalization. BFMs exacerbate this challenge by encoding a vast spectrum of behaviors, ranging from locomotion to multi-contact interactions, which introduces high-dimensional transfer risks. For example, a BFM trained for diverse humanoid motions may fail to adapt to real-world actuator delays, friction variations, or sensor noise, resulting in unstable or unsafe execution. The integration of visual signals into control systems further compounds the challenge, and perceptual domain shifts (e.g., lighting, texture, or camera calibration mismatches) and generalization gaps in visual features can destabilize motion policies based on simulated visual inputs. Current BFMs remain largely confined to simulation, with no documented large-scale real-world deployments. While several pioneering works have been devoted to developing BFM-like controllers in real humanoid robots, such as HugWBC [xue2025unified] and CLONE [li2025clone], their motion skills remain limited, highlighting a significant challenge to Sim2Real feasibility in maintaining behavior richness. This gap stems from behavioral overgeneralization, dynamics mismatches, and latent space instability, which hinder scaling targeted Sim2Real successes (e.g., quadruped locomotion or grasping) to BFM’s complexity.
IV-B2 Computational and System Integration Challenges
While the Sim2Real gap presents deployment challenges at the data and policy levels, another significant challenge in deploying BFMs arises from computational and system integration issues. To ensure responsive motion execution and effective disturbance rejection, the humanoid control policy typically requires real-time command updates at frequencies ranging from 50Hz to 100Hz [li2025bfm, hwangbo2017control]. Unlike simulation environments, where computing resources are plentiful, real-world humanoids operate under strict constraints related to latency, power consumption, and thermal management. Although BFMs can excel with increased model size, as demonstrated by Motivo [tirinzoni2025zeroshot], these constraints limit the feasibility of using larger models. For instance, the BFM-Zero [li2025bfm] only has a maximum parameter count of 44.5M, while SONIC [luo2025sonic] has 42M parameters.
On the other hand, incorporating perceptual input (e.g., RGB or depth images) into BFMs further exacerbates the computational challenge, often requiring additional hardware acceleration to meet real-time processing demands. Devices like the NVIDIA Jetson computing board [biddulph2018comparing, song2025gait] are commonly employed to offload the heavy computational burden of vision-based inference, enabling the robot to process complex multimodal data. However, integrating such hardware introduces significant complexity, particularly in synchronizing the robot’s control system with the inference hardware, managing data communication, and minimizing latency. The increased complexity can complicate overall system coordination, making it difficult to achieve seamless integration between perception, decision-making, and action execution. Future research could address these issues by exploring more efficient sensor fusion methods [alatise2020review], model compression techniques [choudhary2020comprehensive], and hybrid inference architectures [pujol2021fog], etc.
IV-B3 Data Bottleneck